OpenStudy (anonymous):

Streptococcus pneumoniae has a generation time of 30 mins. If you begin your culture at 9am with a single cell, how many cell will be present at noon?

OpenStudy (anonymous):

Define noon, 12:01 pm?

OpenStudy (anonymous):

1200

OpenStudy (anonymous):

It multiplies, at 9:30 there are two. Then at ten they both have split in two so then you have four.

OpenStudy (anonymous):

So what would you have at 12?

OpenStudy (anonymous):

10:30=8 11:00=16 11:30=32 12:00=64

OpenStudy (anonymous):

Hope this is helpful!

OpenStudy (anonymous):

Is everyone orange from ur perspective?

OpenStudy (anonymous):

288?

OpenStudy (nincompoop):

\[ \huge EXPONENTIAL~GROWTH \\ \sf \huge y=y_0e^{kt} \\ where: \\ \large y_0: initial~amount\\ \large k: constant~of~proportionality \\ \large t:time\\ \large y:ending~amount \\ \] EXAMPLE: e. coli is growing in your sink and doubles every 5 minutes. Assuming that it only started with 1 bacterium , how many bacteria could be present after 96 minutes? \( \huge y= y_0 e^{kt} \sf \large find~k~first \\ \huge \color {red}{2} = \color{red}{1}e^{k\times \color{red}{5}} \\ \huge ln 2 = lne^{5k} \\ \huge \frac{ln2}{5} = k \\ \huge .0139 = k ~ \large then ~use~the original~formula \\ \huge y = 1 \times e^{0.139 \times 96} \)